### v2

# MEG: Medical Knowledge-Augmented Large Language Models for Question Answering

Abstract
Question answering is a natural language understanding task that involves reasoning over both explicit context and unstated, relevant domain knowledge. Large language models (LLMs), which underpin most contemporary question answering systems, struggle to induce how concepts relate in specialized domains such as medicine. Existing medical LLMs are also costly to train. In this work, we present MEG, a parameter-efficient approach for medical knowledge-augmented LLMs. MEG uses a lightweight mapping network to integrate graph embeddings into the LLM, enabling it to leverage external knowledge in a cost-effective way. We evaluate our method on four popular medical multiple-choice datasets and show that LLMs greatly benefit from the factual grounding provided by knowledge graph embeddings. MEG attains an average of +10.2% accuracy over the Mistral-Instruct baseline, and +6.7% over specialized models like BioMistral. We also show results based on Llama-3. Finally, we show that MEG’s performance remains robust to the choice of graph encoder.
1 Introduction
Large language models (LLMs) induce knowledge from vast text corpora. Through self-supervised learning, these models capture deeply contextualized representations of input tokens that enable them to generalize to new tasks with remarkable performance. This, as well as their ability to write long coherent passages, has made LLMs incredibly popular, despite their considerable inference costs (Cheng et al., 2023) and their concerning carbon footprint (Strubell et al., 2019). Moreover, current LLMs face significant challenges with handling complex reasoning and ensuring trustworthiness Liu et al. (2023); Huang et al. (2024) and factual consistency Maynez et al. (2020); Zhou et al. (2023); Tam et al. (2023); Hager et al. (2024), essential to critical fields like healthcare. While LLMs are poised to revolutionize our medical system, already performing well on medical licensing exams Jin et al. (2020); Pal et al. (2022); Singhal et al. (2023a); Brin et al. (2023) and other tasks Nazario-Johnson et al. (2023); Van Veen et al. (2023); Tu et al. (2023); Carl et al. (2024), there is still much room for improvement.
To improve reliability and reduce computational costs, researchers have experimented with training from mixtures of corpora and knowledge bases (Pan et al., 2023, 2024). Knowledge Graphs (KGs), such as the Unified Medical Language System (UMLS) Bodenreider (2004), are structured knowledge bases that explicitly store rich factual knowledge. KGs are good at capturing the nuances of complex data and can provide complementary information to LLMs, especially useful for tasks requiring structured understanding. The potential of knowledge-augmented LLMs [^1] In this work, we define a knowledge-augmented LLM as an LLM enhanced with KG embeddings (KGEs). KGEs are dense vector representations of graph entities Ju et al. (2024). Therefore, we also refer to knowledge-augmented LLMs as KGE-augmented LLMs throughout the paper. outlines an interesting research paradigm that can alleviate current challenges of LLMs, and reduce the need of training ever-larger models Hooker (2024). However, how to effectively model interactions between LLMs and KGs remains an open question.
Recent efforts have focused on self-supervised methods for jointly training graph neural networks and pretrained language models Yang et al. (2021); Chien et al. (2022); Brannon et al. (2024). Others Yasunaga et al. (2022); Tang et al. (2024); Plenz and Frank (2024), propose new model architectures to leverage the two modalities, graph and text, during pretraining. These methods learn deep interactions over text and graph, but they require carefully curated pretraining data, are mainly studied for graph-oriented tasks Yang et al. (2021); Chien et al. (2022); Tang et al. (2024), or are yet to be adapted to a generative framework Yasunaga et al. (2022); Plenz and Frank (2024).
In this work, we introduce MEG, a parameter-efficient approach to MEdical knowledGe-augmented LLMs for question answering (QA). We design a lightweight mapping network to unidirectionally translate KG embeddings into the LLM’s vector space. This enables the LLM to interpret the new input embeddings, which, in turn, further conditions its response generation. We use Mistral-Instruct (7B) Jiang et al. (2023) as our base LLM and report results with our best setup: a KG encoder based on GraphSAGE Hamilton et al. (2017) combined with a simple MLP as mapping network. We also provide results with the recently released Llama-3-Instruct (8B) Dubey et al. (2024) as base LLM.
In sum, our contributions are as follows: i) We introduce MEG, a novel approach to knowledge-augmented LLMs based on KGEs. ii) We conduct extensive evaluation on the four popular multiple-choice QA datasets from the MultiMedQA Singhal et al. (2023a) clinical benchmark, and demonstrate the effectiveness of integrating pretrained KGEs into LLMs for medical question answering. Specifically, MEG surpasses strong LLM baselines like BioMistral-7B Labrak et al. (2024) or MediTron-7B Chen et al. (2023), which have followed a costly continued pretraining of the base LLMs on curated biomedical data. iii) We provide insights into the inner workings of MEG, examining the contributions of each module and comparing embedding spaces. We intuitively explain the shifts in the LLM’s representations that drive MEG’s stronger performance. iv) We publicly release the code, trained KGEs and model checkpoints at github.com/lautel/MEG.
2 Related Work
Medical Language Models
Current state-of-the-art (SOTA) in medical QA benchmarks like MedQA Jin et al. (2020), PubMedQA Jin et al. (2019) or MedMCQA Pal et al. (2022) belongs to close-sourced models of unknown size like Med-Gemini Saab et al. (2024), Med-PaLM2 Singhal et al. (2023b) or GPT-4 Nori et al. (2023). Popular open-source LLMs in biomedicine include MedAlpaca Han et al. (2023) and PMC-LLaMA Wu et al. (2023) based on Llama Touvron et al. (2023a), MediTron Chen et al. (2023) based on Llama-2 Touvron et al. (2023b), or BioMistral Labrak et al. (2024) based on Mistral-Instruct Jiang et al. (2023). These models continue pretraining the base general-purpose models on curated medical corpora. More recently, Kim et al. (2024) present the Meerkat models trained with chain-of-thought Wei et al. (2024) synthetic data. Meerkat-7B outperforms the previous best 7B models across several medical benchmarks. However, it takes eight 80G A100 GPUs and 1.5 days to complete training. In contrast, our approach is the first to leverage pretrained medical KGEs and can be trained on four A10G GPUs within a few hours (see § 5 for details).
Knowledge-Augmented Language Models
Bringing together LLMs and KGs is an active line of research that has gained increasing attention from both academia and industry (Pan et al., 2023, 2024). Among numerous efforts in this area, Zhang et al. (2019); Yasunaga et al. (2022); Tang et al. (2024); Zhu et al. (2023), to name a few, propose different methods for combining text and graphs during pretraining. Parallel to these lines of work, Sarmah et al. (2024); Edge et al. (2024); Hu et al. (2024); Mavromatis and Karypis (2024) approach the integration of LLMs and KGs through retrieval-augmented generation (RAG) Lewis et al. (2020). However, the deployment of such knowledge-augmented LLMs for medical QA remains understudied. Our work fills this gap and presents a novel approach to medical knowledge-augmented LLMs based on KGEs. We note that MEG may resemble a sort of RAG system, where an LLM leverages knowledge from an external database of KGEs. In this case, the grounding module would act as retrieval module, fetching appropriate KGEs that ground text information in KG entities as part of a prompt.
3 Problem Formulation
We augment an LLM with KG embeddings to answer medical questions drawing on factual knowledge from the KG. We rely on a large KG in the target domain, namely UMLS Bodenreider (2004). Our proposed approach, MEG, consists of four key components:
  1. i)
    A KG encoder to represent knowledge graph entities in a continuous vector space, while preserving their semantic meaning.
  2. ii)
    An instruction-tuned language decoder capable of generating textual answers.
  3. iii)
    A mapping function $f_{\text{k}}$ that transforms the output of the KG encoder into a representation that can be used by the language decoder. $f_{\text{k}}$ is parameterized by a neural network. Thus, we interchangeably use the term mapping network.
  4. iv)
    A KG grounding module that detects textual entities and grounds them in graph entities.
Figure 1 depicts the full pipeline of MEG. We carefully investigate the design of these components and how they interact with each other (§ 4.1). To attain the best accuracy on downstream tasks, we conduct a two-phase training (§ 4.2).
Definitions.
A generic dataset for multiple-choice question answering (QA) consists of examples with a context paragraph, a question and a candidate answer set, all expressed in text. Given a QA example, each prompt $W$ is the concatenation of context, question and candidate answer set. We denote the sequence of tokens (words) in $W$ as $\{w_{1},\ldots,w_{S}\}$, where $S$ is the maximum sequence length. We denote the sequence of tokens (vectors) in the language model embedding space as $W_{e}=\{w_{e1},\ldots,w_{eS}\}$.
We define a knowledge graph (KG) as a directed graph $G=(V,E)$, where $V$ is the set of entity nodes, and $E\subseteq V\times R\times V$ is the set of edges (triples) that connect nodes in $V$, with $R$ being the set of relation types. Each triple $(s,p,o)$ in a KG represents a knowledge fact, such as ((((Headache, is_a, Cephalgia)))). A KGE $e$ is a mathematical representation that maps each entity $v\in V$ and each relation $r\in R$ of a directed knowledge graph $G$ to low-dimensional vectors in $\mathbb{R}^{g}$, preserving the semantic relationships within the graph.
Finally, we define a KGE-augmented language model to be a function $f_{\text{l}}(W_{e}\oplus f_{\text{k}}(X))$ [^2] Formally, its domain is the set of sequences of elements $x_{i}\in\mathbb{R}^{l}$. with $f_{\text{l}}\in\mathbb{R}^{l}$, where $f_{\text{k}}(X)$ is a set of KGEs, $\{e_{1},\ldots,e_{N}\}$ with $e_{i}\in\mathbb{R}^{g}$, that has been mapped to the LLM’s space using a learned mapping function $f_{k}:\mathbb{R}^{g}\rightarrow\mathbb{R}^{l}$. The language model $f_{\text{l}}$ concatenates these representations to the token word embeddings $W_{e}$ to perform downstream tasks in the fine-tuning steps. A language model is a special case of a KGE-augmented language model with no KGE (N=0).
Figure 1: MEG leverages a pretrained KG encoder and an LLM. During an initial phase of training, MEG learns a mapping network to convert relevant graph features (KGEs) retrieved by the grounding module into token embeddings. During downstream fine-tuning, only the LLM’s weights are updated, keeping the LLM’s embedding layer and mapping network frozen. At inference, the LLM takes the text and the mapped KGEs as input and generates a response.
4 Method
4.1 MEG
MEG combines a pretrained KG encoder and a pretrained LLM by means of an intermediate mapping network (see Figure 1). The KG encoder, which is trained separately on a large medical KG [^3] Specifically, we use UMLS Bodenreider (2004), a widely-used KG in biomedicine with $\sim$300K nodes (entities) and one million edges in total., provides graph embeddings that are directly fed to the mapping network. Then, the LLM uses the text content and mapped KG embeddings as input to generate an answer.
Knowledge Graph Encoder
The KG encoder is trained up-front over the selected graph to generate KGEs. We choose GraphSAGE Hamilton et al. (2017) as our preferred KG encoder. In § 6.1 we present an ablation study with random-walk-based, energy-based translational, and message-passing encoders.
Figure 2: $f_{\text{k}}$ and $g_{\text{k}}$ are embedding transfer functions. $f_{\text{k}}$ takes a set of KGEs $X$ (i.e., node entities) as input, and outputs a mapping of $X$ to the LLM’s vector space. $Y$ is the set of averaged token embeddings of entities in the LLM space. During training, $g_{\text{k}}$ prevents degenerated transformation of graph embeddings. The dashed lines indicate the input for the objective losses.
Mapping Network
The mapping function $f_{\text{k}}$ transforms a sequence of graph features from the KG encoder into a sequence that can be consumed by the LLM. We parameterize $f_{\text{k}}$ as an MLP with four hidden layers of size $d_{h}=128$. In particular, a set of graph embeddings is transformed from $d_{g}=256$ to $d_{l}=4096$ after a series of non-linear transformations through the hidden layers of $f_{\text{k}}$. We denote the initial embedding sets as $X=\{x_{i}\}^{N}_{i=1}$, $Y=\{y_{j}\}^{N}_{j=1}$, $x_{i}\in\mathbb{R}^{d_{g}},y_{j}\in\mathbb{R}^{d_{l}}$, being $x_{i}$ the KGEs, $y_{j}$ the averaged token embeddings of the entity in the LLM, and $N$ the total number of graph embeddings (entities). We further denote the set of mapped embeddings as $f_{\text{k}}(X):=\{f_{\text{k}}(x_{i})\}^{n}_{i=1}$.
The goal is to learn the mapping $f_{\text{k}}$ that transforms $X$ to the LLM’s vector space, while preserving its semantic meaning and structural information. Rather than minimizing the sum of squared differences between $f_{\text{k}}(X)$ and $Y$, we aim at positioning each $x_{i}$ in the neighborhood of its counterpart in $Y$. Pursuing an exact matching of space distributions, such as through a Procrustes transformation Schönemann (1966); Gower (1975), would disregard the structural knowledge encoded in $X$.
To achieve this, we design an architecture similar to Xu et al. (2018) with two mappings $f_{\text{k}}:X\rightarrow Y$ and $g_{\text{k}}:Y\rightarrow X$, as illustrated in Figure 2. We construct an instruction dataset with labels from UMLS’s entities to teach the LLM to interpret the transformed graph embedding $f_{\text{k}}(x_{i})$. Figure 2 shows an example of an instruction, where the placeholder {kg_embedding} is replaced by $f_{\text{k}}(x_{i})$. We train the full network jointly with the LLM [^4] We conducted experiments by training the mapping network and LLM separately on UMLS. However, this approach resulted in worse performance on the downstream tasks tested.. Our loss function consists of three parts: a standard next-token prediction objective (cross-entropy loss $\mathcal{L_{\text{ce}}}$), and a sum of a contrastive loss and a back-translation loss to optimize the mapping network. Specifically,
  • Given a batch $X_{b}$ including a positive pair of examples $x_{i}$ and $x_{j}$, a contrastive objective Hadsell et al. (2006) is a function whose value is low when $x_{i}$ is similar to $x_{j}$ and dissimilar to all others, which are considered negative pairs for $x_{i}$. We employ a popular contrastive self-supervised learning objective Sohn (2016); van den Oord et al. (2019); He et al. (2019), dubbed as NT-Xent loss by Chen et al. (2020). NT-Xent uses dot product as similarity measure, and computes a normalized temperature-scaled cross-entropy loss for a positive pair as follows,
    $\ell_{i,j}=-\log\frac{\exp(x_{i}\cdot x_{j}/\tau)}{\sum_{k=1,k\neq i}^{B}\exp(x_{i}\cdot x_{k}/\tau)},$ (1)
    where B is the batch size and $\tau$ is the temperature. We set the hyper-parameter $\tau~{}=~{}1.0$  [^5] We also evaluate $\tau=0.5$ as in Chen et al. (2020) and $\tau=0.07$ as in He et al. (2019). We choose the final value of $\tau=1.0$ based on accuracy attained on a zero-shot setting on the validation split in MedQA.. The final loss $\mathcal{L_{\text{c}}}$ is computed across all positive pairs in a batch, summed across all batches. Intuitively, the contrastive loss serves as an unsupervised objective function for training the network to bring similar entities closer together in $Y$ and push dissimilar ones apart.
  • We also employ a back-translation loss for preventing degenerated transformation. We enforce that the graph embedding after the forward and the backward transformation should not diverge much from its original direction. Following Xu et al. (2018), we choose the back-translation loss based on cosine similarity. Note that our primary goal is to optimize the forward mapping $f_{\text{k}}:X\rightarrow Y$. Thus, we do not control for back-translation in the reversed path, $g_{\text{k}}:Y\rightarrow X$,
    $\mathcal{L_{\text{bt}}}(f_{\text{k}},g_{\text{k}})=\sum_{i}\left(1-\cos(x_{i},g_{\text{k}}(f_{\text{k}}(x_{i})))\right)$ (2)
Thereby, when training the mapping network jointly with the LLM, we minimize the following objective function:
$\mathcal{L}=\alpha\mathcal{L_{\text{c}}}+\beta\mathcal{L_{\text{bt}}}+\mathcal{L_{\text{ce}}},$ (3)
where $\alpha$ and $\beta$ are scalar hyperparameters to weight each objective in the transformation process. Our network design achieves a good trade-off between expressivity and parameter count, totaling 1.22M parameters. After the mapping is learnt, we freeze the network’s weights and the LLM’s embedding layer during the fine-tuning to downstream tasks (Phase II of training). The backwards transfer network $g_{\text{k}}$ is disconnected and only $f_{\text{k}}$ is used to do the mapping.
Grounding Module
The grounding module takes textual data $W$ as input and links entity mentions in $W$ to their corresponding nodes in the knowledge graph $G$, generating a set of KG embeddings for the LLM to use. A critical step in this process is medical entity disambiguation Vretinaris et al. (2021); Lu et al. (2024), which involves detecting named entities in $W$ and linking them to their unique counterparts in $G$. Since an entity can be referred to in multiple ways, for instance ‘heart attack’ and ‘myocardial infarction’, this step standardizes variations by linking them to a unique identifier in $G$ (following with the previous example, both mentions correspond to Concept Unique Identifier (CUI) ‘C0155626’ in UMLS). This grounding ensures retrieval of relevant information for each example. We use the entity linker presented in Neumann et al. (2019) [^6] We use the last version of scispaCy (v2.5.0), which supports linking to UMLS and has near 3M unique concepts., which covers 99% of the concepts mentioned in the MedMentions dataset Mohan and Li (2019) and 86% of the concepts mentioned in the MedQA dataset Jin et al. (2020). These two datasets provide ground truth UMLS annotations.
4.2 Training
We aim to train MEG to achieve competent results on medical question answering benchmarks while minimizing computational cost. To do this, we conceive a two-phase training strategy with a minimal part of the model’s parameters updated.
Phase I: Embedding Transfer Learning
We first learn the optimal transformation $f_{\text{k}}:X\rightarrow Y$ so that the mapped KG embeddings retain relevant information from the KG and can be effectively used by the LLM. As explained in § 4.1, we create an instruction dataset to guide the LLM in learning the relationship between its original representation of medical entities and their mapped graph embeddings. The train set contains 297,927 examples, following the same template shown in Figure 2 for every entity label in UMLS. [^7] We investigate whether data augmentation at this stage could lead to more accurate results in downstream tasks. We augment the initial $\sim$300K examples by creating new instructions with multiple entities, e.g.‘Explain to me these medical concepts: […]’, to better match the setting from downstream tasks, which often include several entities per sample. This process doubles the dataset size, with an augmentation that normally distributes the number of entities per instruction between 2 and 10. Results are within ±$\pm 0.2$ accuracy in MedQA compared to training without the augmented data. Due to the extra computational costs and minor (if any) gains, we did not explore this option further. We train for one epoch jointly the mapping network and the LLM to minimize the objective from Eq. 3.
Phase II: Downstream task
Given a medical multiple choice QA dataset, we fine-tune MEG to answer the input question based on the textual content and information leveraged from the mapped KG embeddings. We format the input prompts $W$ as follows. For each example in the dataset, we concatenate the context (if any), question and candidate answer set following the pseudo-code shown in Figure 3. The placeholder {kg_embedding} is replaced with $N$ KGEs transformed by the mapping network888 In both training phases, we investigate the effect of injecting the mapped KGEs at the last layer of the LLM instead of after the embedding layer. These early experiments revealed little to no impact on zero-shot downstream accuracy, but slightly worsen the fine-tuned accuracy as measured on the validation set of MedQA with three random initialization seeds. This finding suggests that the LLM benefits from attending the external KGEs during fine-tuning, enabling more contextualized representations of these embeddings.. We assume $f_{\text{k}}$ is learned in phase I, so we keep the mapping network weights frozen and disconnect the backward network $g_{\text{k}}$, which merely served to regulate the learning of $f_{\text{k}}$ and prevent degenerated transformation. Similarly, the LLM’s embedding layer is also frozen.
[INST] Please address the following medical question based on the Input text and any useful information you may find in the given concepts from a medical graph. Input: context question Options: {for option in options %} {{letter}}) {{text}} {endfor %} Answer with the best option directly. Ignore irrelevant information. Graph: {{kg_embeddings}} [/INST] Answer: {{correct_option}}
Figure 3: Template used to generate instructions for all QA datasets. The context is optional, depending on the dataset. At inference time, the text after [/INST] is generated by the language model.
5 Experimental Details
Data
Following previous research on medical LLMs, we evaluate MEG on four well-known medical benchmarks that require extensive background knowledge. The first one is MedQA-USMLE (MedQA) Jin et al. (2020), which consists of 10,178 train questions and 1,273 test questions, formatted with four choices each. The content was originally curated by experts from the US Medical License Exam. The second benchmark, PubMedQA Jin et al. (2019), was collected from PubMed abstracts and includes 1,000 expert labeled question-answer pairs. The task is to produce a yes/no/maybe answer based on the question and an abstract as context. As previously done by others Singhal et al. (2023a); Chen et al. (2023); Labrak et al. (2024), we use 500 random [^9] We split the data following a similar distribution of answers between train and test splits. samples for evaluation. The remaining 500 samples, though limited in size, serve as our only source of training data. We exclude the 211k artificially labeled yes/no samples provided by Jin et al. (2019) to avoid bias towards these two options. The third benchmark, MedMCQA Pal et al. (2022), contains 179,722101010We detect 3,100 duplicate questions in the train split, which we remove. train questions from Indian medical entrance exams. Due to the unavailability of answer keys for the test set, we follow others Wu et al. (2023); Tu et al. (2023); Labrak et al. (2024) and report results on the validation set (4,183 questions). Lastly, MMLU-Medical Singhal et al. (2023a) includes 1,089 questions, each with four options, across six medical and biology-related categories drawn from Hendrycks et al. (2021). Since this dataset only provides test data, we evaluate the generalization performance of MEG fine-tuned on MedMCQA as in Chen et al. (2023). Thus, results on MMLU-Medical report out-of-distribution inference.
Training Details
We initialize the KG node embeddings with token embeddings from SapBERT Liu et al. (2020). SapBERT leverages contextualized embeddings from a pretrained BERT-based language model for biomedical KGs like UMLS. This initialization leads to improved performance compared to random embedding initialization. We train GraphSAGE with same hyperparameters as in Hamilton et al. (2017).
During phase I of training, described in § 4.2, we randomly initialize the mapping network and load the pretrained weights of the LLM. We fully train the mapping network and perform low-rank adaptation (LoRA, Hu et al. (2022)) fine-tuning on every linear layer of the LLM, while the remaining parameters are frozen. This parameter-efficient tuning approach allows to learn the equivalent of 2% of the model’s parameters. Our full architecture results on 216M trainable parameters for MEG-Mistral. After fine-tuning, we merge the LLM’s updated parameters with the base model. Since the input prompt has a fixed size (see Figure 2), we use a reduced sequence length (124) to optimize computational efficiency. We train for one epoch with gradient accumulation over 8 steps to achieve an effective batch size of 128. We employ a cosine learning rate scheduler with learning rate of $51e-$ 5, warmup ratio of 3% and no weight decay. We use mixed precision (bfloat16) and FlashAttention2 Dao (2023) to optimize memory usage and speed up computations on the LLM. Training takes 4h on 4 NVIDIA A10G GPUS using DeepSpeed [^11] https://www.deepspeed.ai/ for distributed training.
After phase I, the models are evaluated in a zero-shot setup on downstream tasks, already able to use information from graph embeddings. To fine-tune on a specific task, we perform phase II of training. To allow batching, the number of KGEs injected to the LLM is fixed across samples121212 The average number of ground entities per instance varies across datasets according to the median number of ground entities. We set $N=20$ in MedQA, PubMedQA and ‘professional medicine’ in MMLU-Medical; $N=3$ in MedMCQA and $N=2$ in the remaining categories from MMLU-Medical.. If the grounding module retrieves more KGEs, we randomly select $N$. Otherwise, we add zero-padding. We assume the mapping network is learned in phase I, so we freeze its weights and disconnect the backward network $g_{\text{k}}$. We also freeze the LLM’s embedding layer. This approach reduces the computational complexity and speeds up fine-tuning with a total of 83M trainable parameters in both MEG-Mistral and MEG-Llama. Fine-tuning is done for 3 epochs, with a sequence length of 400 (500 for PubMedQA), learning rate of $41e-$ 4 and effective batch size of 32. The remaining hyperparameters have the same value as in phase I. We did not optimize hyperparameters for Llama-3-Instruct.
6 Results
We evaluate accuracy on four medical multiple-choice question datasets in three variants of MEG: MEG-Mistral1 and MEG-Mistral3, based on Mistral-7B-Instruct-v0.1 and -v0.3, respectively; and MEG-Llama, based on Llama-3-Instruct. We report average accuracy and standard deviation across three random seeds. Our results in Tables 1 and 2 reveal consistent average improvement across datasets compared to baselines.
In-prompt graph triples provide useful information
We investigate whether the inclusion of KG information can positively influence the LLM’s answers. To establish a primary baseline, we take Mistral-Instruct-v0.1 and MedQA as a running example. For each question, we select a maximum of 10 named entities $s$ and randomly retrieve 2 graph neighbors $o$ for each, resulting in a maximum of 20 graph triples $(s,p,o)$. We include them as part of the prompt, in natural language [^13] We append the triples at the end of the instruction in JSONL-style, i.e., [{s1,p1,o1}, {s2,p2,o2}, …].. Table 1 shows a degradation in zero-shot accuracy when including triples to the prompt. This can be due to the random selection of the final triples (to fit in the context length), since the semantic information varies among them significantly. This limitation speaks in favor of using KGEs to condense representation of entities to a single embedding. Also, as Hager et al. (2024) point out, LLMs’ face difficulties in interpreting large amounts of information. However, fine-tuning the model with triples boosts accuracy, even surpassing BioMistral, a model adapted from Mistral-Instruct-v0.1 through continued pretraining on curated biomedical data. This baseline highlights the value of graph data for the LLM, but it still does not fully leverage the structural and semantic information provided by the KG.
Model Acc
ZS Mistral-Instruct-v0.1 42.3±0.3
BioMistral 44.4±0.2
Mistral-Instruct-v0.1 w/ graph 40.4±0.4
FT Mistral-Instruct-v0.1 42.0±0.2
BioMistral 50.6±0.3
Mistral-Instruct-v0.1 w/ graph 52.7±0.2
Table 1: Ablation study on the utility of the explicit information encoded in knowledge graph triples. We report accuracy on MedQA. ZS stands for ‘zero-shot’; FT stands for ‘fine-tuning’. Results from Labrak et al. (2024).
MedQA PubMedQA MedMCQA MMLU-Medical Avg MMLU-Medical
Human (pass) 60.0 60.0 Clinical K. Genetics Anatomy P. Medicine C. Biology C. Medicine
Human (expert) 87.0 78.0 90.0
Models based on Llama Models based on Llama
MedAlpaca (7B) 40.1±0.4 73.6±0.3 37.0±0.3 55.1±1.1 51.4 53.1±0.9 58.0±2.2 54.1±1.6 58.8±0.3 58.1±1.3 48.6±0.5
MEDITRON-7B∗‡ 52.0±- 74.4±- 59.2±- 54.2±- 60.0 57.2±- 64.6±- 49.3±- 55.4±- 53.8±- 44.8±-
Meerkat-8B∗‡ 74.2±- - 62.7±- 75.2±- 70.7 74.3±- 76.7±- 74.8±- 75.3±- 76.1±- 74.3±-
MEG-Llama (8B) 66.0±0.2 78.0±0.3 60.6±0.3 74.9±0.7 69.9 72.3±0.5 83.0±1.5 64.5±0.7 79.4±0.3 80.6±0.4 69.4±0.9
Models based on Mistral-Instruct 7B Models based on Mistral-Instruct 7B
Mistral-Instruct-v0.1 42.0±0.2 73.8±${}^{\triangledown}_{\pm{0.4}}$ 46.1±0.1 59.1±1.0 55.3 62.9±0.2 57.0±0.8 55.6±1.0 59.4±0.6 62.5±1.0 57.2±2.1
BioMistral 50.6±0.3 77.5±0.1 48.1±0.2 59.1±1.3 58.8 59.9±1.2 64.0±1.6 56.5±1.8 60.4±0.5 59.0±1.5 54.7±1.0
BioMistral DARE 51.1±0.3 77.7±0.1 48.7±0.1 61.9±1.2 59.9 62.3±1.3 67.0±1.6 55.8±0.9 61.4±0.3 66.9±2.3 58.0±0.5
Meerkat-7B 70.3±- - 60.6±- 70.5±- 67.1 71.6±- 74.8±- 63.2±- 77.3±- 70.8±- 65.2±-
MEG-Mistral1 54.6±0.2 74.6±0.6 56.4±0.4 60.3±0.9 61.5 58.1±0.8 68.7±0.2 54.4±0.5 62.9±0.9 61.1±2.2 56.6±1.0
MEG-Mistral3 60.8±0.2 74.4±0.5 58.4±0.6 68.2±0.4 65.5 64.9±0.2 69.6±0.8 63.0±1.0 72.8±0.4 73.6±0.0 65.2±0.2
Table 2: Main results on four medical multiple-choice question answering benchmarks (left) and fine-grained results on MML-Medical (right). We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Avg stands for average across datasets. Results reproduced by Labrak et al. (2024). Reproduced by us with the same data splits used in this work. Results from the original papers.
KGE-augmented LLMs show accuracy gains
To fully exploit the KG’s rich structural information, we replace the text triples by node embeddings. This approach also compacts the KG’s information into a much shorter input sequence. Table 2 shows performance on the four medical multiple-choice question answering datasets (left), and a fine-grained evaluation on MMLU-Medical subjects (right). Across all datasets, MEG-Mistral consistently outperforms the baselines’ accuracy, and the merged model DARE, the best BioMistral model in Labrak et al. (2024). The only exception is PubMedQA, where BioMistral surpasses MEG. Their higher accuracy may result from using the artificially labeled training set. Instead, we rather train with the small subset of manually labeled samples to avoid biasing the model towards yes/no answers (see § 5). The high accuracy on MMLU-Medical indicates that MEG retains good generalization capabilities. Current SOTA for 7B models, Meerkat-7B, proves the effectiveness of training on chain-of-thought (CoT) in-domain synthetic data. Future work includes exploring CoT instruction tuning in our phase I of training, exploiting information from graph triples instead of relying only on entity labels.
6.1 Ablation study
In this section, we evaluate how the choice of graph encoder and mapping network architecture impact MEG-Mistral1’s performance on a downstream task (case study on MedQA).
On the impact of the graph encoder
We train encoders based on random-walk (RDF2Vec Ristoski and Paulheim (2016)), energy (DistMult Yang et al. (2015)) and message-passing (GraphSAGE Hamilton et al. (2017) and eGraphSAGE, an edge-type-aware variant inspired by Hu et al. (2020)’s adaptation). Along with their impact in MEG’s performance, we include a link classification task as a proxy to evaluate their capabilities. Since these encoders are fundamentally distinct, they capture diverse graph properties, as reflected in classification accuracy in Figure 4, plain (orange) bars. eGraphSAGE stands out with a considerably higher score (73.9), as it naturally integrates edge-type information during training.
However, higher accuracy in a graph-oriented task such as link classification, does not lead to better performance in a language-oriented downstream task in MEG. When we integrate these KGEs in MEG-Mistral1 and evaluate zero-shot and fine-tune settings on MedQA (Figure 4, stripped bars), eGraphSAGE’s substantial advantage in link classification does not carry over to MEG-Mistral1, as evidenced by the smaller performance gap across encoders (ranging from 52.1 to 54.2). This suggests that the role of the KGEs in our setup aligns with our intuition: they guide the answer generation by activating the LLM’s semantic region that leads to the correct answer. The difference is more notable in a zero-shot setting, where RDF2Vec produces the highest rate of not valid answers (NA). DistMult’s lower NA rate indicates it may better align with the LLM’s embedding space.
Figure 4: Ablation study on KG encoder choice. Plain bars show edge classification accuracy over UMLS; stripped bars show MEG-Mistral1’s zero-shot (//) and fine-tuned (\\) accuracy on MedQA; the dashed line represents accuracy with random embeddings; red dots mark the ratio of not valid answers (NA) in the zero-shot setting.
Parameters (M) MedQA (ZS) MedQA (FT)
MLP 2$\times$512 4.98 41.6 53.5
MLP 2$\times$384 3.74 41.5 53.3
Transf. 4$\times$128 2.18 39.9 53.2
MLP 4$\times$128 1.22 41.6 54.2
MLP 3$\times$128 1.18 42.7 52.7
MLP 2$\times$128 1.15 38.2 51.3
Table 3: Comparison of MEG-Mistral1’s accuracy on MedQA with different mapping networks. ZS: zero-shot; FT: fine-tuned.
On the impact of the mapping network
To assess the impact of the mapping network architecture we replace our 4-layer, 128-dimensional MLP (MLP 4$\times$128) with the alternative designs from Table 3. Our final choice, MLP 4$\times$128, outperforms all others in the fine-tuning setting while maintaining a small size. The transformer’s lower score highlights the unnecessary overhead of attention layers, as the mapping network’s task of embedding transformation does not benefit from attending the input sequence.
6.2 Qualitative Analysis
This section provides insights into the representation spaces before and after the mapping network, comparing them with the LLM’s vector space.
Visualizing the embeddings
To track the mapped KGEs, we select three UMLS concepts (entities) representing different semantic and specificity levels within the graph’s hierarchy, measured by the number of descendants ( is_a and subclass_of relations): ‘Diabetes Mellitus’ (CUI: C0011849, a broad disease category), ‘Headache’ (CUI: C0018681, a symptom), and ‘Atorvastatin Calcium’ (CUI: C0286650, a specific pharmacological substance).
Figure 5: t-SNE visualization of the embeddings: before and after the mapping network. After mapping, the relative KGEs’ structure along hierarchy levels is preserved, albeit slightly rotated (e.g., in Level 4, see the diagonal gap which hints the orientation of the blobs) and with reversed sparsity. Note the clustering effect over contextualized label embeddings: the mapped KGEs draw them to a specific region.
Figure 5 depicts t-SNE [^14] We have also witnessed the same overall behavior using UMAP and MDS to visualize the embeddings. plots of the concepts (Level 0) and their hierarchies (Level 1 to  4) in the KGE space (top row) and the mapped KGE along with the concept’s contextualized embeddings in the LLM space (bottom row). Such contextualized embeddings are the LLM’s token embeddings generated for the labels of the concepts when verbalizing their KG’s triples, i.e., they represent the concepts in all the contexts given by the KG.
Examining the distribution of concepts in the upper and lower rows, we observe two effects of the mapping. First, the relative structure of KGEs is preserved after mapping to the LLM’s space, albeit slightly rotated (see Level 1 and Level 4) and with reversed sparsity (high-density groups become less dense, while low-density groups tighten; yet the global between-group structure is the same) [^15] Given that KGE’s and LLM’s embeddings represent two spaces with different dimensionalities, we applied t-SNE separately to each; thus, a difference in scale is expected.. Second, the clustering effect of the contextualized label embeddings (see Level 0): while the labels (circles in the bottom row) are scattered, the mapped embeddings cover such spaces. As we descend in the graph’s hierarchy, the contextualized labels become increasingly scattered, yet the mapped KGEs, while remaining within the same region, stay more cohesive. Note that these contextualized labels are distinct vectors for the same label across different contexts. Indeed, their sparsity suggests that they do not completely capture the conceptual dimension of the entity they represent. The mapping seems to activate the LLM’s embedding region associated with each concept, providing semantic information the model uses to condition its answer generation.
Probing MEG for acquired knowledge
We informally probe MEG to estimate its acquired medical knowledge. We handcraft a set of open-ended questions to query Mistral-Instruct before and after the phase I training on UMLS (§ 4.2), and MEG. As an example, we show an excerpt (irrelevant information to the question is omitted for brevity) of models’ answers to the question ‘What are the active and inactive compounds present in Atorvastatin?’ [^16] This question targets Atorvastatin (C0286650 in UMLS). and highlight the key differences. MEG uses the KGE for Atorvastatin and provides a more knowledgeable response.
Mistral-Instruct (baseline): …The active compound in Atorvastatin is Atorvastatin itself, while the inactive compounds are Atorvastatin calcium salt and Atorvastatin magnesium salt.
Mistral-Instruct after phase I: …The active compound present in Atorvastatin is Atorvastatin calcium. …is responsible for the drug’s therapeutic effect. There are no inactive compounds present in Atorvastatin.
MEG-Mistral1: …the active compound Atorvastatin calcium.The inactive compounds present in Atorvastatin include excipients such as microcrystalline cellulose, croscarmellose sodium, magnesium stearate, and sodium lauryl sulfate
7 Analysis and Discussion
Our approach is efficient not only because it trains a small fraction of the LLM’s parameters but also because it seamlessly handles out-of-vocabulary terms. A new term’s KGE can be initialized, for instance, by averaging its one-hop neighbors in the KG, making MEG both lightweight and adaptable to new vocabulary. The efficacy of this method should be evaluated in future work.
Garikipati et al. (2024) demonstrate that prompt engineering can outperform fine-tuning in medical QA for open-source LLMs. However, our focus was to investigate the viability of integrating knowledge from KG embeddings into LLMs rather than optimizing for peak downstream performance. Our experiments show that supervised fine-tuning of KGE-augmented LLMs yields more accurate answers than other specialized baselines. Chain-of-thought tuning, as shown by Kim et al. (2024), is another promising step forward to improve MEG’s accuracy. MEG improves response generation by injecting KGEs in a single generation step. This suggests that MEG could also benefit from chain-of-thought tuning, as each of the reasoning steps would increase precision of the model’s response.
Besides, the sensitivity of LLMs to the information order in multiple-choice questions, also known as positional bias, is well-documented Pezeshkpour and Hruschka (2023); Zheng et al. (2024). More specific to biomedicine, Liévin et al. (2024) and Hager et al. (2024) show how variations in sequence can significantly impact diagnostic accuracy of medical-aligned language models. However, the robustness to changes in information order remain understudied in most medical model evaluations. Recent studies Wang et al. (2024b, a) find that greedy decoding combined with text answer evaluation gives more consistent answers compared to first-token evaluation, particularly for instruction-tuned LLMs. Thus, in an attempt to alleviate this issue, we inspect the text answer generated by the model instead of ranking the candidate answers by the log probability of its first token prediction. However, whether KGEs further help the LLMs in mitigating positional biases—as well as other Lyell and Coiera (2017); Moor et al. (2023); Ness et al. (2024) biases— needs to be explored in future work.
8 Conclusion
We introduce MEG, a novel medical knowledge-augmented LLM based on KGEs for question answering tasks. To the best of our knowledge, we are the first to inject pretrained KGEs into an LLM via a lightweight mapping network, enabling the model to interpret structural graph information from the medical domain. We present a comprehensive evaluation on four medical multiple-choice question benchmarks, revealing that LLMs can highly benefit from the factual information encoded in KG embeddings. Our results suggest that integrating KGEs with LLMs offers a promising path towards specialized language models.
Acknowledgments
We are grateful to CoAStaL members for their comments on earlier versions of this project. Special thanks to Rita Ramos, Constanza Fierro and Jonas F. Lotz for their insightful feedback and to Nicolas Garneau for his support during initial experiments. Laura Cabello was supported by the Novo Nordisk Foundation (grant NNF 20SA0066568). The work of Carlos Bobed was supported by the I+D+i project PID2020-113903RB-I00 (funded by MCIN/AEI/ 10.13039/501100011033) and the project T42_23R (funded by Gobierno de Aragón, Spain).