### v3

# MEG: Medical Knowledge-Augmented Large Language Models for Question Answering

Abstract
Question answering is a natural language understanding task that involves reasoning over both explicit context, and unstated relevant domain knowledge. Despite the high cost of training, large language models (LLMs)—the backbone of most modern question-answering systems—still struggle to reliably capture the nuanced relationships between concepts that are crucial for reasoning in specialized fields like medicine. In this work, we present MEG, a parameter-efficient approach for medical knowledge-augmented LLMs. MEG uses a lightweight mapping network to incorporate knowledge graph embeddings into the LLM, enabling it to leverage external knowledge in a cost-effective way. We evaluate our method on four popular medical multiple-choice datasets and show that LLMs i) can effectively interpret knowledge graph embeddings and ii) gain significant advantages from the factual grounding these embeddings provide. MEG attains an average of +6.7% and +9.9% accuracy over specialized models like BioMistral-7B and MediTron-7B, respectively. Finally, we show that MEG’s performance remains robust to the choice of graph encoder.
1 Introduction
Large language models (LLMs) induce knowledge from vast text corpora. Through self-supervised learning, these models capture deeply contextualized representations of input tokens that enable them to generalize to new tasks with remarkable performance. This, as well as their ability to write long coherent passages, has made LLMs incredibly popular, despite their considerable inference costs (Cheng et al., 2023) and their concerning carbon footprint (Strubell et al., 2019). Moreover, current LLMs face significant challenges with handling complex reasoning and ensuring trustworthiness Liu et al. (2023); Huang et al. (2024) and factual consistency Maynez et al. (2020); Zhou et al. (2023); Tam et al. (2023); Hager et al. (2024), essential to critical fields like healthcare. While LLMs are poised to revolutionize our medical system, already performing well on medical licensing exams Jin et al. (2020); Pal et al. (2022); Singhal et al. (2023a); Brin et al. (2023) and other tasks Nazario-Johnson et al. (2023); Van Veen et al. (2023); Tu et al. (2023); Carl et al. (2024), there is still much room for improvement.
To improve reliability and reduce computational costs, researchers have experimented with training from mixtures of corpora and knowledge bases (Pan et al., 2023, 2024). Knowledge Graphs (KGs), such as the Unified Medical Language System (UMLS) Bodenreider (2004), are structured knowledge bases that explicitly store rich factual knowledge. KGs are good at capturing the nuances of complex data and can provide complementary information to LLMs, especially useful for tasks requiring structured understanding. However, this complementary information is introduced as a new modality: graphs. Therefore, it extends beyond the plain text that LLMs can naturally interpret, requiring additional mechanisms to process and integrate it effectively. The potential of knowledge-augmented LLMs [^1] In this work, we define a knowledge-augmented LLM as an LLM enhanced with KG embeddings (KGEs). KGEs are dense vector representations of graph entities Ju et al. (2024). Therefore, we also refer to knowledge-augmented LLMs as KGE-augmented LLMs throughout the paper. outlines an interesting research paradigm that can alleviate current challenges of LLMs, and reduce the need of training ever-larger models Hooker (2024). However, how to effectively model interactions between LLMs and KGs remains an open question.
Recent efforts have focused on self-supervised methods for jointly training graph neural networks and pretrained language models Yang et al. (2021); Chien et al. (2022); Brannon et al. (2024). Others Yasunaga et al. (2022); Tang et al. (2024); Plenz and Frank (2024), propose new model architectures to leverage the two modalities, graph and text, during pretraining. These methods learn deep interactions over text and graph, but they require carefully curated pretraining data, are mainly studied for graph-oriented tasks Yang et al. (2021); Chien et al. (2022); Tang et al. (2024), or are yet to be adapted to a generative framework Yasunaga et al. (2022); Plenz and Frank (2024).
In this work, we introduce MEG, a parameter-efficient approach to MEdical knowledGe-augmented LLMs for question answering (QA). MEG eliminates the need to train base LLMs for specialized domains by incorporating relevant KGEs at inference time. To achieve this, we design a lightweight mapping network that unidirectionally translates KGEs into the LLM’s vector space, allowing the model to interpret these embeddings, which, in turn, further conditions its response generation. Notably, KGEs are injected directly at token level, after the model’s embedding layer. We experiment with Mistral-Instruct (7B) Jiang et al. (2023) and Llama-3-Instruct (8B) Dubey et al. (2024) as our base models, and report results with our best setup for KGEs: a KG encoder based on GraphSAGE Hamilton et al. (2017) combined with a simple Multilayer Perceptron (MLP) as mapping network.
In sum, our contributions are as follows: [^2] Code and KGEs will be published upon acceptance. i) We introduce MEG, a novel approach to knowledge-augmented LLMs based on KGEs [^3] While we frame our approach in the medical domain, we would like to remark that it is domain independent given that it is built on any large KG. However, as we focused in this domain given the abundance of previous results and datasets, we refrain from making further generalized claims (in spite of there being strong evidences supporting them).. MEG is an efficient approach to domain specialization of models, as it updates only a small fraction of the LLM’s parameters while providing access to a large, reliable knowledge source (a KG). ii) We conduct extensive evaluation on the four popular multiple-choice QA datasets from the MultiMedQA Singhal et al. (2023a) clinical benchmark, and demonstrate the effectiveness of integrating pretrained KGEs into LLMs for medical question answering. Specifically, MEG surpasses well-establish biomedical LLM baselines like BioMistral-7B Labrak et al. (2024) or MediTron-7B Chen et al. (2023), which have followed a costly continued pretraining of the base LLMs on curated biomedical data. iii) We provide insights into MEG’s dual modality by examining the contributions of each module and comparing embedding spaces of KGEs and LLMs. We intuitively explain the shifts in the LLM’s representations that drive MEG’s stronger performance.
2 Related Work
Medical Language Models
Current state-of-the-art (SOTA) in medical QA benchmarks like MedQA Jin et al. (2020), PubMedQA Jin et al. (2019) or MedMCQA Pal et al. (2022) belongs to close-sourced models of unknown size like Med-Gemini Saab et al. (2024), Med-PaLM2 Singhal et al. (2023b) or GPT-4 Nori et al. (2023). Popular open-source LLMs in biomedicine include MedAlpaca Han et al. (2023) and PMC-LLaMA Wu et al. (2023) based on Llama Touvron et al. (2023a), MediTron Chen et al. (2023) based on Llama-2 Touvron et al. (2023b), or BioMistral Labrak et al. (2024) based on Mistral-Instruct Jiang et al. (2023). These models continue pretraining the base general-purpose models on curated medical corpora. More recently, Kim et al. (2024) present the Meerkat models trained with chain-of-thought Wei et al. (2024) synthetic data. Meerkat-7B outperforms the previous best 7B models across several medical benchmarks. However, it takes eight 80G A100 GPUs and 1.5 days to complete training. In contrast, our approach is among the first to leverage pretrained medical KGEs to condition the LLM’s generation and can be trained on four A10G GPUs within a few hours (see § 5 for details).
Knowledge-Augmented Language Models
Bringing together LLMs and KGs is an active line of research that has gained increasing attention from both academia and industry (Pan et al., 2023, 2024). Among numerous efforts in this area, Zhang et al. (2019); Yasunaga et al. (2022); Tang et al. (2024); Zhu et al. (2023), to name a few, propose different methods for combining text and graphs during pretraining. Parallel to these lines of work, Sarmah et al. (2024); Edge et al. (2024); Hu et al. (2024); Mavromatis and Karypis (2024) approach the integration of LLMs and KGs through retrieval-augmented generation (RAG) Lewis et al. (2020). However, the deployment of such knowledge-augmented LLMs for medical QA remains understudied. Our work fills this gap and presents a novel approach to medical knowledge-augmented LLMs based on KGEs. We note that MEG may resemble a sort of RAG system, where an LLM leverages knowledge from an external database of KGEs. However, our approach does not retrieve document chunks to include them in the textual prompt. Instead, MEG fetches KGEs of named entities from the user’s prompt and injects these embeddings into the LLM, enriching the query’s context and guiding answer generation. For a comprehensive benchmark of RAG in medical QA, we refer the reader to Xiong et al. (2024).
3 Problem Formulation
Our objective is to augment a general purpose LLM with specialized knowledge, so it can generate accurate answers without the need of fine-tuning to the new domain. We propose to use a large KG as the source of specialized knowledge. The reason for this is two-folded: a KG is a reliable source of factual knowledge and it captures structured relationship between concepts, that can potentially augment the LLM’s representation of such concepts. This design choice marks the multimodality of our solution: text generation will be conditioned on embedded representations of graph concepts.
Thus, our main task is to augment an LLM with KG embeddings to answer questions in a relevant domain. Our proposed approach, MEG, consists of four key components:
  1. i)
    A KG encoder to represent knowledge graph entities in a continuous vector space, while preserving their semantic meaning.
  2. ii)
    An instruction-tuned language decoder capable of generating textual answers.
  3. iii)
    A mapping function $f_{\text{k}}$ that transforms the output of the KG encoder into a representation that can be used by the language decoder. $f_{\text{k}}$ is parameterized by a neural network. Thus, we interchangeably use the term mapping network.
  4. iv)
    A KG grounding module that detects textual entities and grounds them in graph entities.
Figure 1 depicts our full pipeline. We carefully investigate the design of these components and how they interact with each other (§ 4.1). The mapping network is jointly trained with the LLM to learn an appropriate transformation of KGE, and it is frozen in downstream applications (§ 4.2).
Definitions.
A generic dataset for multiple-choice question answering (QA) consists of examples with a context paragraph, a question and a candidate answer set, all expressed in text. Given a QA example, each prompt $W$ is the concatenation of context, question and candidate answer set. We denote the sequence of tokens (words) in $W$ as $\{w_{1},\ldots,w_{S}\}$, where $S$ is the maximum sequence length. We denote the sequence of tokens (vectors) in the language model embedding space as $W_{e}=\{w_{e1},\ldots,w_{eS}\}$.
We define a knowledge graph (KG) as a directed graph $G=(V,E)$, where $V$ is the set of entity nodes, and $E\subseteq V\times R\times V$ is the set of edges (triples) that connect nodes in $V$, with $R$ being the set of relation types. Each triple $(s,p,o)$ in a KG represents a knowledge fact, such as ((((Headache, is_a, Cephalgia)))). A KGE $e$ is a mathematical representation that maps each entity $v\in V$ and each relation $r\in R$ of a directed knowledge graph $G$ to low-dimensional vectors in $\mathbb{R}^{g}$, preserving the semantic relationships within the graph.
Finally, we define a KGE-augmented language model to be a function $f_{\text{l}}(W_{e}\oplus f_{\text{k}}(X))$ [^4] Formally, its domain is the set of sequences of elements $x_{i}\in\mathbb{R}^{l}$. with $f_{\text{l}}\in\mathbb{R}^{l}$, where $f_{\text{k}}(X)$ is a set of KGEs, $\{e_{1},\ldots,e_{N}\}$ with $e_{i}\in\mathbb{R}^{g}$, that has been mapped to the LLM’s space using a learned mapping function $f_{k}:\mathbb{R}^{g}\rightarrow\mathbb{R}^{l}$. The language model $f_{\text{l}}$ concatenates these representations to the token word embeddings $W_{e}$ to perform downstream tasks in the fine-tuning steps. A language model is a special case of a KGE-augmented language model with no KGE (N=0).
Figure 1: MEG leverages a pretrained KG encoder and an LLM. During an initial phase of training, MEG learns a mapping network to convert relevant graph features (KGEs) retrieved by the grounding module into token embeddings. During downstream fine-tuning, only a subset of the LLM’s weights is updated, while the embedding layer and mapping network remain frozen. At inference, the LLM takes the text and the mapped KGEs as input and generates a response.
4 Method
4.1 MEG
MEG combines a pretrained KG encoder and a pretrained LLM by means of an intermediate mapping network (see Figure 1). The KG encoder, which is trained separately on a large medical KG [^5] Specifically, we use UMLS Bodenreider (2004), a widely-used KG in biomedicine with $\sim$300K nodes (entities) and one million edges in total., provides graph embeddings that are directly fed to the mapping network. The mapping network outputs embeddings that are injected into the LLM after its embedding layer, influencing the LLM’s answer generation.
Knowledge Graph Encoder
The KG encoder is trained up-front over the selected graph to generate KGEs. We choose GraphSAGE Hamilton et al. (2017) as our preferred KG encoder. In § 6.1 we present an ablation study with random-walk-based, energy-based translational, and message-passing encoders.
Figure 2: $f_{\text{k}}$ and $g_{\text{k}}$ are embedding transfer functions. $f_{\text{k}}$ takes a set of KGEs $X$ (i.e., node entities) as input, and outputs a mapping of $X$ to the LLM’s vector space. $Y$ is the set of averaged token embeddings of entities in the LLM space. During training, $g_{\text{k}}$ prevents degenerated transformation of graph embeddings. The dashed lines indicate the input for the objective losses.
Mapping Network
The mapping function $f_{\text{k}}$ transforms a sequence of graph features from the KG encoder into a sequence that can be consumed by the LLM. We parameterize $f_{\text{k}}$ as an MLP with four hidden layers of size $d_{h}=128$. In particular, a set of graph embeddings is transformed from $d_{g}=256$ to $d_{l}=4096$ after a series of non-linear transformations through the hidden layers of $f_{\text{k}}$. We denote the initial embedding sets as $X=\{x_{i}\}^{N}_{i=1}$, $Y=\{y_{j}\}^{N}_{j=1}$, $x_{i}\in\mathbb{R}^{d_{g}},y_{j}\in\mathbb{R}^{d_{l}}$, being $x_{i}$ the KGEs, $y_{j}$ the averaged token embeddings of the entity in the LLM, and $N$ the total number of graph embeddings (entities). We further denote the set of mapped embeddings as $f_{\text{k}}(X):=\{f_{\text{k}}(x_{i})\}^{n}_{i=1}$.
The goal is to learn the mapping $f_{\text{k}}$ that transforms $X$ to the LLM’s vector space, while preserving its semantic meaning and structural information. Rather than minimizing the sum of squared differences between $f_{\text{k}}(X)$ and $Y$, we aim at positioning each $x_{i}$ in the neighborhood of its counterpart in $Y$. Pursuing an exact match of space distributions, such as through a Procrustes transformation Schönemann (1966); Gower (1975), would disregard the structural knowledge encoded in $X$.
To achieve this, we design an architecture similar to Xu et al. (2018) with two mappings $f_{\text{k}}:X\rightarrow Y$ and $g_{\text{k}}:Y\rightarrow X$, as illustrated in Figure 2. We construct an instruction dataset with labels from UMLS’s entities to teach the LLM to interpret the transformed graph embedding $f_{\text{k}}(x_{i})$. Figure 2 (top) shows an example of an instruction, where the placeholder {kg_embedding} is replaced by $f_{\text{k}}(x_{i})$ at token level. We train the full network jointly with the LLM [^6] We conducted experiments by training the mapping network and LLM separately on UMLS. However, this approach resulted in worse performance on the downstream tasks tested.. Our loss function consists of three parts: a standard next-token prediction objective (cross-entropy loss $\mathcal{L_{\text{ce}}}$), and a sum of a contrastive loss and a back-translation loss to optimize the mapping network. Specifically,
  • Given a batch $X_{b}$ including a positive pair of examples $x_{i}$ and $x_{j}$, a contrastive objective Hadsell et al. (2006) is a function whose value is low when $x_{i}$ is similar to $x_{j}$ and dissimilar to all others, which are considered negative pairs for $x_{i}$. We employ a popular contrastive self-supervised learning objective Sohn (2016); van den Oord et al. (2019); He et al. (2019), dubbed as NT-Xent loss by Chen et al. (2020). NT-Xent uses dot product as similarity measure, and computes a normalized temperature-scaled cross-entropy loss for a positive pair as follows,
    $\ell_{i,j}=-\log\frac{\exp(x_{i}\cdot x_{j}/\tau)}{\sum_{k=1,k\neq i}^{B}\exp(x_{i}\cdot x_{k}/\tau)},$ (1)
    where B is the batch size and $\tau$ is the temperature. We set the hyper-parameter $\tau~{}=~{}1.0$  [^7] We also evaluate $\tau=0.5$ as in Chen et al. (2020) and $\tau=0.07$ as in He et al. (2019). We choose the final value of $\tau=1.0$ based on accuracy attained on a zero-shot setting on the validation split in MedQA.. The final loss $\mathcal{L_{\text{c}}}$ is computed across all positive pairs in a batch, summed across all batches. Intuitively, the contrastive loss serves as an unsupervised objective function for training the network to bring similar entities closer together in $Y$ and push dissimilar ones apart.
  • We also employ a back-translation loss for preventing degenerated transformation. We enforce that the graph embedding after the forward and the backward transformation should not diverge much from its original direction. Following Xu et al. (2018), we choose the back-translation loss based on cosine similarity. Note that our primary goal is to optimize the forward mapping $f_{\text{k}}:X\rightarrow Y$. Thus, we do not control for back-translation in the reversed path, $g_{\text{k}}:Y\rightarrow X$,
    $\mathcal{L_{\text{bt}}}(f_{\text{k}},g_{\text{k}})=\sum_{i}\left(1-\cos(x_{i},g_{\text{k}}(f_{\text{k}}(x_{i})))\right)$ (2)
Thereby, when training the mapping network jointly with the LLM, we minimize the following objective function:
$\mathcal{L}=\alpha\mathcal{L_{\text{c}}}+\beta\mathcal{L_{\text{bt}}}+\mathcal{L_{\text{ce}}},$ (3)
where $\alpha$ and $\beta$ are scalar hyperparameters to weight each objective in the transformation process. Our network design achieves a good trade-off between expressivity and parameter count, totaling 1.22M parameters. After the mapping is learnt, we freeze the network’s weights and the LLM’s embedding layer during the fine-tuning to downstream tasks (Phase II of training). The backwards transfer network $g_{\text{k}}$ is disconnected and only $f_{\text{k}}$ is used to do the mapping.
Grounding Module
The grounding module takes textual data $W$ as input and links mentioned entities in $W$ to their corresponding nodes in the knowledge graph $G$, generating the corresponding entity mentions’ KGEs for the LLM to use. These KG embeddings will enrich the question’s context, helping to guide the answer. A critical step in this process is medical entity disambiguation Vretinaris et al. (2021); Lu et al. (2024), which involves detecting named entities in $W$ and linking them to their unique counterparts in $G$. Since an entity can be referred to in multiple ways, for instance ‘heart attack’ and ‘myocardial infarction’, this step standardizes variations by linking them to a unique identifier in $G$ (following with the previous example, both mentions correspond to Concept Unique Identifier (CUI) ‘C0155626’ in UMLS). This grounding ensures retrieval of relevant information for each example. We use the entity linker presented in Neumann et al. (2019) [^8] We use the last version of scispaCy (v2.5.0), which supports linking to UMLS and has near 3M unique concepts., which covers 99% of the concepts mentioned in the MedMentions dataset Mohan and Li (2019) and 86% of the concepts mentioned in the MedQA dataset Jin et al. (2020). These two datasets provide ground truth UMLS annotations.
4.2 Training
We aim to train MEG to achieve competent results on medical question answering benchmarks while minimizing computational cost. To do this, We perform an initial training phase for embedding transfer learning.
Phase I: Embedding Transfer Learning
We first learn the optimal transformation $f_{\text{k}}:X\rightarrow Y$ so that the mapped KG embeddings retain relevant information from the KG and can be effectively used by the LLM. As explained in § 4.1, we create an instruction dataset from UMLS labels to guide the LLM in learning the relationship between its original representation of medical entities and their mapped graph embeddings. The train set contains 297,927 examples, following the same template shown in Figure 2 for every entity label in UMLS. [^9] We investigate whether data augmentation at this stage could lead to more accurate results in downstream tasks. We augment the initial $\sim$300K examples by creating new instructions with multiple entities, e.g.‘Explain to me these medical concepts: […]’, to better match the setting from downstream tasks, which often include several entities per sample. This process doubles the dataset size, with an augmentation that normally distributes the number of entities per instruction between 2 and 10. Results are within ±$\pm 0.2$ accuracy in MedQA compared to training without the augmented data. Due to the extra computational costs and minor (if any) gains, we did not explore this option further. We train for one epoch jointly the mapping network and the LLM to minimize the objective from Eq. 3. At this stage, MEG can leverage concepts from the KG to augment its domain knowledge and generate better, informed answers.
Fine-tuning on Downstream Tasks
Given a medical multiple choice QA dataset, we fine-tune MEG to answer the input question based on the textual content and information leveraged from the mapped KG embeddings. We format the input prompts $W$ as follows. For each example in the dataset, we concatenate the context (if any), question and candidate answer set following the pseudo-code shown in Figure 3.
At the token level, the placeholder embedding {kg_embedding} is replaced with $N$ transformed KGEs produced by the mapping network101010 In both training phases, we investigate the effect of injecting the mapped KGEs at the last layer of the LLM instead of after the embedding layer. These early experiments revealed little to no impact on zero-shot downstream accuracy, but slightly worsen the fine-tuned accuracy as measured on the validation set of MedQA with three random initialization seeds. This finding suggests that the LLM benefits from attending the external KGEs during fine-tuning, enabling more contextualized representations of these embeddings.. The mapping network weights remain frozen, and the backward network $g_{\text{k}}$ is effectively disabled, which previously served to regulate the learning of $f_{\text{k}}$ and prevent degenerated transformations. Similarly, the LLM’s embedding layer is also frozen. This approach reduces the computational complexity and speeds up fine-tuning.
[INST] Please address the following medical question based on the Input text and any useful information you may find in the given concepts from a medical graph. Input: context question Options: {for option in options %} {{letter}}) {{text}} {endfor %} Answer with the best option directly. Ignore irrelevant information. Graph: {{kg_embeddings}} [/INST] Answer: {{correct_option}}
Figure 3: Template used to generate instructions for all QA datasets. The context is optional, depending on the dataset. At inference time, the text after [/INST] is generated by the language model.
5 Experimental Details
Data
Following previous research on medical LLMs, we evaluate MEG on four well-known medical benchmarks that require extensive background knowledge. The first one is MedQA-USMLE (MedQA) Jin et al. (2020), which consists of 10,178 train questions and 1,273 test questions, formatted with four choices each. The content was originally curated by experts from the US Medical License Exam. The second benchmark, PubMedQA Jin et al. (2019), was collected from PubMed abstracts and includes 1,000 expert labeled question-answer pairs. The task is to produce a yes/no/maybe answer based on the question and an abstract as context. As previously done by others Singhal et al. (2023a); Chen et al. (2023); Labrak et al. (2024), we use 500 random [^11] We split the data following a similar distribution of answers between train and test splits. samples for evaluation. The remaining 500 samples, though limited in size, serve as our only source of training data. We exclude the 211k artificially labeled yes/no samples provided by Jin et al. (2019) to avoid bias towards these two options. The third benchmark, MedMCQA Pal et al. (2022), contains 179,722121212We detect 3,100 duplicate questions in the train split, which we remove. train questions from Indian medical entrance exams. Due to the unavailability of answer keys for the test set, we follow others Wu et al. (2023); Tu et al. (2023); Labrak et al. (2024) and report results on the validation set (4,183 questions). Lastly, MMLU-Medical Singhal et al. (2023a) includes 1,089 questions, each with four options, across six medical and biology-related categories drawn from Hendrycks et al. (2021). Since this dataset only provides test data, we evaluate the generalization performance of MEG fine-tuned on MedMCQA as in Chen et al. (2023). Thus, results on MMLU-Medical report out-of-distribution inference.
Training Details
We initialize the KG node embeddings with token embeddings from SapBERT Liu et al. (2020). SapBERT leverages contextualized embeddings from a pretrained BERT-based language model for biomedical KGs like UMLS. This initialization leads to improved performance compared to random embedding initialization. We train GraphSAGE with same hyperparameters as in Hamilton et al. (2017).
During phase I of training, described in § 4.2, we randomly initialize the mapping network and load the pretrained weights of the LLM. We fully train the mapping network and perform low-rank adaptation (LoRA, Hu et al. (2022)) fine-tuning on every linear layer of the LLM (except for the embedding layer), while the remaining parameters are frozen. This parameter-efficient tuning approach allows to learn the equivalent of 2% of the model’s parameters. Our full architecture results on 216M trainable parameters.After training, we merge the LLM’s updated parameters with the base model. Training takes 4h on 4 NVIDIA A10G GPUS using DeepSpeed [^13] https://www.deepspeed.ai/ for distributed training. See Appendix A for further details.
After phase I, the models are evaluated in a zero-shot setup on downstream tasks, already able to use information from graph embeddings. To achieve optimal performance in downstream tasks, we fine-tune the models as in § 4.2. To allow batching, the number of KGEs injected to the LLM is fixed across samples141414 The average number of ground entities per instance varies across datasets according to the median number of ground entities. We set $N=20$ in MedQA, PubMedQA and ‘professional medicine’ in MMLU-Medical; $N=3$ in MedMCQA and $N=2$ in the remaining categories from MMLU-Medical.. If the grounding module retrieves more KGEs, we randomly select $N$. Otherwise, we add zero-padding. See Appendix A for training details.
Model Acc
ZS Mistral-Instruct-v0.1 42.3±0.3
BioMistral 44.4±0.2
Mistral-Instruct-v0.1 w/ graph 40.4±0.4
FT Mistral-Instruct-v0.1 42.0±0.2
BioMistral 50.6±0.3
Mistral-Instruct-v0.1 w/ graph 52.7±0.2
Table 1: Ablation study on the utility of the information encoded in knowledge graph triples expressed in natural language. We report accuracy on MedQA. ZS stands for ‘zero-shot’; FT stands for ‘fine-tuning’. Results from Labrak et al. (2024).
6 Results
We evaluate accuracy on four medical multiple-choice question datasets in three variants of MEG: MEG-Mistral1 and MEG-Mistral3, based on Mistral-7B-Instruct-v0.1 and -v0.3, respectively; and MEG-Llama, based on Llama 3 Instruct (8B). We report average accuracy and standard deviation across three random seeds. Our results in Tables 1 and 2 reveal consistent average improvement across datasets compared to baselines. We adopted both LLMs to show that the benefits of our proposal come regardless the underlying model we took as starting point.
In-prompt graph triples provide useful information
We investigate whether the inclusion of KG information can positively influence the LLM’s answers. To establish a primary baseline, we take Mistral-Instruct-v0.1 and MedQA as a running example. For each question, we select a maximum of 10 named entities $s$ and randomly retrieve 2 graph neighbors $o$ for each, resulting in a maximum of 20 graph triples $(s,p,o)$. We include them as part of the prompt, in natural language [^15] We append the triples at the end of the instruction in JSONL-style, i.e., [{s1,p1,o1}, {s2,p2,o2}, …].. Table 1 shows a degradation in zero-shot accuracy when including triples to the prompt. This can be due to the random selection of the final triples (to fit in the context length), since the semantic information varies among them significantly. This limitation speaks in favor of using KGEs to condense representation of entities to a single embedding. Also, as Hager et al. (2024) point out, LLMs’ face difficulties in interpreting large amounts of information. However, fine-tuning the model with triples boosts accuracy, even surpassing BioMistral, a model adapted from Mistral-Instruct-v0.1 through continued pretraining on curated biomedical data. This baseline highlights the value of graph data for the LLM, but it still does not fully leverage the structural and semantic information provided by the KG.
MedQA PubMedQA MedMCQA MMLU-Medical Avg
Models based on Llama
Llama 3 Instruct (8B) 60.3±1.4 75.8±1.4 56.4±0.8 72.8±3.3 66.3
MedAlpaca (7B) 40.1±0.4 73.6±0.3 37.0±0.3 55.1±1.1 51.4
MEDITRON (7B) 52.0±- 74.4±- 59.2±- 54.2±- 60.0
Meerkat-8B 74.2±- - 62.7±- 75.2±- 70.7
RAG-MEDITRON (70B) 49.6±1.4 56.4±2.2 52.7±0.8 65.4±1.4 56.0
RAG-Llama 2 (70B) 44.9±1.4 50.4±2.2 43.1±0.8 54.6±1.5 48.3
MEG-Llama (8B) 66.0±0.2 78.0±0.3 60.6±0.3 74.9±0.7 69.9
Models based on Mistral 7B
Mistral-Instruct-v0.1 42.0±0.2 73.8±${}^{\triangledown}_{\pm{0.4}}$ 46.1±0.1 59.1±1.0 55.3
BioMistral 50.6±0.3 77.5±0.1 48.1±0.2 59.1±1.3 58.8
BioMistral DARE 51.1±0.3 77.7±0.1 48.7±0.1 61.9±1.2 59.9
Meerkat-7B 70.3±- - 60.6±- 70.5±- 67.1
RAG-Mixtral (8$\times$7B) 60.0±1.4 67.6±2.1 56.4±0.8 75.9±1.3 65.0
MEG-Mistral1 54.6±0.2 74.6±0.6 56.4±0.4 60.3±0.9 61.5
MEG-Mistral3 60.8±0.2 74.4±0.5 58.4±0.6 68.2±0.4 65.5
Table 2: Main results on four medical multiple-choice question answering benchmarks. We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Avg stands for average across datasets. Llama 3 Instruct and Mistral-Instruct are our base, general purpose models. Models with gray background are baselines fine-tuned on biomedical data. MEG variants outperformed the respective baselines. Results reproduced by Labrak et al. (2024). Results from Xiong et al. (2024). Reproduced by us with the same data splits used in this work. Results from the original papers.
KGE-augmented LLMs show accuracy gains
To fully exploit the KG’s rich structural information, we use node embeddings to guide the LLM’s generation. This approach also compacts the KG’s information into a much shorter input sequence. Table 2 reports accuracy on four multiple-choice question datasets. Across all datasets, MEG-Llama and both MEG-Mistral consistently outperform the base LLMs and the fine-tuned baselines (in gray). The only exception is PubMedQA, where BioMistral surpasses both MEG-Mistral. Its higher accuracy may result from using the artificially labeled training set. Instead, we rather train with the small subset of manually labeled samples to avoid biasing the model towards yes/no answers (see § 5). The high accuracy on MMLU-Medical indicates that MEG retains good generalization capabilities. See Appendix B for a fine-grained evaluation on MMLU-Medical subjects. Current SOTA for 7B models, Meerkat-7B, proves the effectiveness of training on chain-of-thought (CoT) in-domain synthetic data. Future work includes exploring CoT instruction tuning in our phase I of training, exploiting information from graph triples instead of relying only on entity labels. For reference, we also include RAG results from Xiong et al. (2024). Despite models used by Xiong et al. (2024) are larger than ours and we cannot directly compare to them, accuracy results indicate that MEG is a solid, lightweight alternative.
6.1 Ablation study
In this section, we take MEG-Mistral1 as reference and evaluate the impact of the choice of graph encoder and mapping network architecture on a downstream task (case study on MedQA).
On the impact of the graph encoder
We train encoders based on random-walk (RDF2Vec Ristoski and Paulheim (2016)), energy (DistMult Yang et al. (2015)) and message-passing (GraphSAGE Hamilton et al. (2017) and eGraphSAGE, an edge-type-aware variant inspired by Hu et al. (2020)’s adaptation). Along with their impact in MEG’s performance, we include a link classification task as a proxy to evaluate their capabilities. Since these encoders are fundamentally distinct, they capture diverse graph properties, as reflected in classification accuracy in Figure 4, plain (orange) bars. eGraphSAGE stands out with a considerably higher score (73.9), as it naturally integrates edge-type information during training.
However, higher accuracy in a graph-oriented task such as link classification, does not lead to better performance in a language-oriented downstream task in MEG. When we integrate these KGEs in MEG-Mistral1 and evaluate zero-shot and fine-tune settings on MedQA (Figure 4, stripped bars), eGraphSAGE’s substantial advantage in link classification does not carry over to MEG-Mistral1, as evidenced by the smaller performance gap across encoders (ranging from 52.1 to 54.2). This suggests that the role of the KGEs in our setup aligns with our intuition: they guide the answer generation by activating the LLM’s semantic region that leads to the correct answer. The difference is more notable in a zero-shot setting, where RDF2Vec produces the highest rate of not valid answers (NA). DistMult’s lower NA rate indicates it may better align with the LLM’s embedding space.
Figure 4: Ablation study on KG encoder choice. Plain bars show edge classification accuracy over UMLS; stripped bars show MEG-Mistral1’s zero-shot (//) and fine-tuned (\\) accuracy on MedQA; the dashed line represents accuracy with random embeddings; red dots mark the ratio of not valid answers (NA) in the zero-shot setting.
Parameters (M) MedQA (ZS) MedQA (FT)
MLP 2$\times$512 4.98 41.6 53.5
MLP 2$\times$384 3.74 41.5 53.3
Transf. 4$\times$128 2.18 39.9 53.2
MLP 4$\times$128 1.22 41.6 54.2
MLP 3$\times$128 1.18 42.7 52.7
MLP 2$\times$128 1.15 38.2 51.3
Table 3: Ablation study on the mapping network done on a separated 1K subset of the train set. Results show accuracy on MedQA for MEG-Mistral1. ZS: zero-shot; FT: fine-tuned.
On the impact of the mapping network
To assess the impact of the mapping network architecture we replace our 4-layer, 128-dimensional MLP (MLP 4$\times$128) with the alternative designs from Table 3. Our final choice, MLP 4$\times$128, outperforms all others in the fine-tuning setting while maintaining a small size. The transformer’s lower score highlights the unnecessary overhead of attention layers, as the mapping network’s task of embedding transformation does not benefit from attending the input sequence.
6.2 Qualitative Analysis
This section provides insights into the representation spaces before and after the mapping network, comparing them with the LLM’s vector space.
Visualizing the embeddings
To track the mapped KGEs, we select three UMLS concepts (entities) representing different semantic and specificity levels within the graph’s hierarchy, measured by the number of descendants ( is_a and subclass_of relations): ‘Diabetes Mellitus’ (CUI: C0011849, a broad disease category), ‘Headache’ (CUI: C0018681, a symptom), and ‘Atorvastatin Calcium’ (CUI: C0286650, a specific pharmacological substance).
Figure 5: t-SNE visualization of the embeddings: before and after the mapping network. After mapping, the relative KGEs’ structure along hierarchy levels is preserved, albeit slightly rotated (e.g., in Level 4, see the diagonal gap which hints the orientation of the blobs) and with reversed sparsity. Note the clustering effect over contextualized label embeddings: the mapped KGEs draw them to a specific region.
Figure 5 depicts t-SNE [^16] We have also witnessed the same overall behavior using UMAP and MDS to visualize the embeddings. plots of the concepts (Level 0) and their hierarchies (Level 1 to  4) in the KGE space (top row) and the mapped KGE along with the concept’s contextualized embeddings in the LLM space (bottom row). Such contextualized embeddings are the LLM’s token embeddings generated for the labels of the concepts when verbalizing their KG’s triples, i.e., they represent the concepts in all the contexts given by the KG.
Examining the distribution of concepts in the upper and lower rows, we observe two effects of the mapping. First, the relative structure of KGEs is preserved after mapping to the LLM’s space, albeit slightly rotated (see Level 1 and Level 4) and with reversed sparsity (high-density groups become less dense, while low-density groups tighten; yet the global between-group structure is the same) [^17] Given that KGE’s and LLM’s embeddings represent two spaces with different dimensionalities, we applied t-SNE separately to each; thus, a difference in scale is expected.. Second, the clustering effect of the contextualized label embeddings (see Level 0): while the labels (circles in the bottom row) are scattered, the mapped embeddings cover such spaces. As we descend in the graph’s hierarchy, the contextualized labels become increasingly scattered, yet the mapped KGEs, while remaining within the same region, stay more cohesive. Note that these contextualized labels are distinct vectors for the same label across different contexts. Indeed, their sparsity suggests that they do not completely capture the conceptual dimension of the entity they represent. The mapping seems to activate the LLM’s embedding region associated with each concept, providing semantic information the model uses to condition its answer generation.
Probing MEG for acquired knowledge
We informally probe MEG to estimate its acquired medical knowledge. We handcraft a set of open-ended questions to query Mistral-Instruct before and after the phase I training on UMLS (§ 4.2), and MEG-Mistral1. As an example, we show an excerpt (irrelevant information to the question is omitted for brevity) of models’ answers to the question ‘What are the active and inactive compounds present in Atorvastatin?’ [^18] This question targets Atorvastatin (C0286650 in UMLS). and highlight the key differences. MEG uses the KGE for Atorvastatin and provides a more knowledgeable response.
Mistral-Instruct (baseline): …The active compound in Atorvastatin is Atorvastatin itself, while the inactive compounds are Atorvastatin calcium salt and Atorvastatin magnesium salt.
Mistral-Instruct after phase I: …The active compound present in Atorvastatin is Atorvastatin calcium. …is responsible for the drug’s therapeutic effect. There are no inactive compounds present in Atorvastatin.
MEG-Mistral1: …the active compound Atorvastatin calcium.The inactive compounds present in Atorvastatin include excipients such as microcrystalline cellulose, croscarmellose sodium, magnesium stearate, and sodium lauryl sulfate
7 Discussion and Future Research Lines
Our approach is efficient not only because it trains a small fraction of the LLM’s parameters but also because it seamlessly handles out-of-vocabulary terms. We present and efficient approach to domain specialization of models, as it updates only a small fraction of the LLM’s parameters while seamlessly handling out-of-vocabulary terms. A new term’s KGE can be initialized, for instance, by averaging its one-hop neighbors in the KG, making MEG both lightweight and adaptable to new vocabulary. However, the efficacy of this method should be evaluated in future work.
Garikipati et al. (2024) demonstrate that prompt engineering can outperform fine-tuning in medical QA for open-source LLMs. However, our focus was to investigate the viability of integrating knowledge from KG embeddings into LLMs rather than optimizing for peak downstream performance. Our experiments show that supervised fine-tuning of KGE-augmented LLMs yields more accurate answers than other specialized baselines. Chain-of-thought tuning, as shown by Kim et al. (2024), is another promising step forward to improve MEG’s accuracy. MEG improves response generation by injecting KGEs in a single generation step. This suggests that MEG could also benefit from chain-of-thought tuning, as each of the reasoning steps would increase precision of the model’s response.
Besides, the sensitivity of LLMs to the information order in multiple-choice questions, also known as positional bias, is well-documented Pezeshkpour and Hruschka (2023); Zheng et al. (2024). More specific to biomedicine, Liévin et al. (2024) and Hager et al. (2024) show how variations in sequence can significantly impact diagnostic accuracy of medical-aligned language models. However, the robustness to changes in information order remain understudied in most medical model evaluations. Recent studies Wang et al. (2024b, a) find that greedy decoding combined with text answer evaluation gives more consistent answers compared to first-token evaluation, particularly for instruction-tuned LLMs. Thus, in an attempt to alleviate this issue, we inspect the text answer generated by the model instead of ranking the candidate answers by the log probability of its first token prediction. However, whether KGEs further help the LLMs in mitigating positional biases—as well as other Lyell and Coiera (2017); Moor et al. (2023); Ness et al. (2024) biases—needs to be explored in future work.
8 Conclusion
We introduce MEG, a novel medical knowledge-augmented LLM based on KGEs for question answering tasks. To the best of our knowledge, we are the first to inject pretrained KGEs into an LLM via a lightweight mapping network. The LLM interprets the graph information encoded in KGEs, which further guides the response generation. We present a comprehensive evaluation on four medical multiple-choice question benchmarks, revealing that LLMs can highly benefit from the factual information encoded in KG embeddings. Our results suggest that integrating KGEs with LLMs offers a promising path towards specialized language models.
Appendix A Training Details
A.1 Phase I: Embedding Transfer Learning
Since the input prompt has a fixed size (see Figure 2), we use a reduced sequence length of 124 to optimize computational efficiency. We train for one epoch with gradient accumulation over 8 steps to achieve an effective batch size of 128. We employ a cosine learning rate scheduler with learning rate of $51e-$ 5, warmup ratio of 3% and no weight decay. We use mixed precision (bfloat16) and FlashAttention2 Dao (2023) to optimize memory usage and speed up computations on the LLM.
A.2 Fine-tuning on Downstream Tasks
Fine-tuning is done for 3 epochs, with a sequence length of 400 (500 for PubMedQA), learning rate of $41e-$ 4 and effective batch size of 32. The remaining hyperparameters have the same value as in Phase I. We did not optimize hyperparameters for Llama-3-Instruct. Total number of trainable parameters: 83M, in both MEG-Mistral and MEG-Llama.
Appendix B Extended Results
Clinical K. Genetics Anatomy P. Medicine C. Biology C. Medicine
Models based on Llama
LLama 3 Instruct (8B) 73.6±2.7 85.0±3.6 68.9±4.0 69.1±2.8 78.5±3.4 61.8±3.7
MedAlpaca (7B) 53.1±0.9 58.0±2.2 54.1±1.6 58.8±0.3 58.1±1.3 48.6±0.5
MEDITRON-7B∗‡ 57.2±- 64.6±- 49.3±- 55.4±- 53.8±- 44.8±-
Meerkat-8B∗‡ 74.3±- 76.7±- 74.8±- 75.3±- 76.1±- 74.3±-
MEG-Llama (8B) 72.3±0.5 83.0±1.5 64.5±0.7 79.4±0.3 80.6±0.4 69.4±0.9
Models based on Mistral 7B
Mistral-Instruct-v0.1 62.9±0.2 57.0±0.8 55.6±1.0 59.4±0.6 62.5±1.0 57.2±2.1
BioMistral 59.9±1.2 64.0±1.6 56.5±1.8 60.4±0.5 59.0±1.5 54.7±1.0
BioMistral DARE 62.3±1.3 67.0±1.6 55.8±0.9 61.4±0.3 66.9±2.3 58.0±0.5
Meerkat-7B 71.6±- 74.8±- 63.2±- 77.3±- 70.8±- 65.2±-
MEG-Mistral1 58.1±0.8 68.7±0.2 54.4±0.5 62.9±0.9 61.1±2.2 56.6±1.0
MEG-Mistral3 64.9±0.2 69.6±0.8 63.0±1.0 72.8±0.4 73.6±0.0 65.2±0.2
Table 4: Performance on six individual MMLU-Medical categories. We report accuracy ($\uparrow$) and standard deviation ($\downarrow$), when available, of other 7B and 8B medical open-source models. Models with gray background are baselines fine-tuned on biomedical data.